OpenStudy (chasebraves):

A purse contains 21 coins consisting of nickels and dimes. Find the number of coins of each kind if the total value is $1.65. If 5n represents the value of the nickels, then which of the following expressions represents the value of the dimes?

rebeccaxhawaii (rebeccaxhawaii):

1.65=5n+10d 21=n+d

OpenStudy (agl202):

\[n+d=21\] \[10d+5n=170\] \[From (1),d=21-n\] Substitute into 2 and solve for n then d.

rebeccaxhawaii (rebeccaxhawaii):

its not 170 its 1.65 @Agl202

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (agl202):

Thx for correcting me my bad